Find out the value of an investment that will give you Tk. 60,000 per year forever with a

discount rate of 14% compounded quarterly.


1
Expert's answer
2021-12-07T11:21:27-0500

A= P(1 + r/n)n.t

60000 = P(1 + 0.14/4)4

60000 = P(1+ 0.035)4

60000 = P(1.148)

60000 = 1.148P

P = 52286.53

= Tk. 52286.53